If the energy required to change the state of ice to water was 1.68MJ and the specific latent heat of fusion of ice is 226000J/Kg. If this same liquid water has a specific heat capacity of 4200J/Kg degrees celsius and was heated so it's temperature rose by 15 degrees celsius, what extra energy must be supplied to do this.
1
Expert's answer
2019-01-16T10:21:24-0500

The mass of ice

m=Q_1/λ=(1.68×〖10〗^6 J)/(226 000 J/kg)=7.43 kg

So, the extra energy

Q_2=cmΔT=4200 J/kg×7.43 kg×15℃=0.468×〖10〗^6 J=0.468 MJ